OpenStudy (anonymous):

Simplify the expression. 3[(15-3)^2 divide 4] A; 9 B; 108 C; 18 D; 36

OpenStudy (anonymous):

\[3[(15-3)^2\div4\]

zepdrix (zepdrix):

\[\large 3\left[(\color{red}{15-3})^2 \div 4\right]\]Understand what to do first? Let's deal with the red part, simplify the inner brackets before we deal with the exponent.

zepdrix (zepdrix):

\[\large 3\left[(\color{red}{12})^2 \div 4\right]\]That part make sense?

OpenStudy (anonymous):

Yeah 15-3 = 12.

zepdrix (zepdrix):

\[\large 3\left[\color{red}{(\color{black}{12})^2} \div 4\right]\]Ok good, now we want to deal with the exponent. :) What does 12^2 give you?

OpenStudy (anonymous):

36

zepdrix (zepdrix):

\(\large 12^2\) is the same as \(\large 12 \times 12\).

OpenStudy (anonymous):

Right.

zepdrix (zepdrix):

It's a bit larger than 36 :) If you can't remember, maybe use a calculator for this step.

zepdrix (zepdrix):

Oh or did you mean 36 for the final answer? :D

OpenStudy (anonymous):

Oh whoops I did 112x12 divided by 4 hold on.

OpenStudy (anonymous):

12x12 divided by 4*

OpenStudy (anonymous):

12x12 = 144.

zepdrix (zepdrix):

Yah that sounds right, if you jumped ahead and did the divide by 4, yes you get 36.

OpenStudy (anonymous):

That was actually pretty easy.

zepdrix (zepdrix):

So that leaves you with this right?\[\large 3\left[36\right]\]

OpenStudy (anonymous):

Yeah. So the correct answer for this problem is 36?

zepdrix (zepdrix):

No we have a 3 being multiplied on the outside. See the brackets?

zepdrix (zepdrix):

3 x 36 Don't let the fact that they're square brackets confuse you. Round brackets and square brackets mean the same thing, multiplication.

OpenStudy (anonymous):

Right.

zepdrix (zepdrix):

So we get a final answer of 3 x 36 = 108

OpenStudy (anonymous):

Ohh. I see now.

zepdrix (zepdrix):

The things that's tricky about a problem like this is that you need to remember your `order of operations`. PEMDAS P - Parentheses first E - Exponents next M/D - Then multiplication and division, from left to right A/S - Then finally, addition and subtraction, from left to right.

OpenStudy (anonymous):

Copied that.

zepdrix (zepdrix):

cool c:

OpenStudy (anonymous):

Thanks a lot man!

zepdrix (zepdrix):

np c: